OpenStudy (anonymous):

Regarding linear differential equations, why is the integrating factor \[e ^{\int\P(x)}\], where did that come from?

OpenStudy (anonymous):

It comes from the product rule.

myininaya (myininaya):

\[y'+p(x)y=q(x)\] we need to multiply both sides by v(x) such that v*p=v' so that we can write vy'+vpy=vq remember we chose a v such that vp=v' vy'+v'y=vq (yv)'=vq and so on... any ways what is v well v'=dv/dx so \[\frac{dv}{dx}=vp\] we can use separation of variables to find v \[\frac{1}{v} dv=p dx\] integrate both sides \[lnv=\int\limits_{}^{}pdx+C\] \[v=e^{\int\limits_{}^{}p d}+C\]

myininaya (myininaya):

v>0

myininaya (myininaya):

any questions?

myininaya (myininaya):

by the way we don't need the constant we can take the constant to be zero

myininaya (myininaya):

but it doesn't matter you can always divide the constant out in the equation if you do side to multiply by it

OpenStudy (anonymous):

Excellent, Thank you.